1). Given the following vectors: u=(1, 2, -4) and v=(2, 3, 5)

Find:

a) u.v

b) u.u

c) (u.v) v

d) angle between u and v

e) unit vector


2. )Suppose that B =(2, 2, 1) and makes 30° with A and A•B = 6. What is the magnitude of A?


3. )What is the angle between A and B if A• B = 0


4.) Find (u+ v)•(2u – v) given that u•u= 4 u•v =-5 and v•v=10


5. )Determine if u and u = (cos 0, sin 0,–1); v=(sin 0,-cos 0,0) V are orthogonal


1
Expert's answer
2021-11-15T16:44:36-0500

1)

a)

"u\cdot v=2+2\cdot 3-4\cdot 5=-12"


b)

"u\cdot u=1+2\cdot 2+4\cdot 4=21"


c)

"(u\cdot v) v=-12(2, 3, 5)=(-24,-36,-60)"


d)

"u\cdot v=|u||v|cos\alpha"

"|u|=\sqrt{1+2^2+4^2}=\sqrt{21}"

"|v|=\sqrt{2^2+3^2+5^2}=\sqrt{38}"


"cos\alpha=-\frac{12}{\sqrt{21\cdot38}}=-0.4248"

"\alpha=arccos(-0.4248)=115\degree"


e)

"\tilde{u}=u/|u|=(1/\sqrt{21}, 2/\sqrt{21}, -4/\sqrt{21})"

"\tilde{v}=v/|v|=(2/\sqrt{38}, 3/\sqrt{38}, 5/\sqrt{38})"


2)

"A•B =|A||B|cos30\degree= 6"

"|B|=\sqrt{2^2+2^2+1}=3"


"|A|=\frac{A\cdot B}{|B|cos30\degree}=\frac{6}{3\sqrt 3/2}=4/\sqrt 3"


3)

"A•B =|A||B|cos\alpha= 0"

"\alpha=90\degree"


4)

"(u+ v)•(2u – v)=2 u•u+u\cdot v- v•v=2\cdot4-5-10=-7"


5)

"u\cdot v = (cos 0, sin 0,–1)\cdot(sin 0,-cos 0,0)=0+0+0=0"

u and v are are orthogonal